Solve the differential equation of the following homogeneous equation.


xydx + (x^2 + y^2)dy = 0


1
Expert's answer
2021-09-20T10:52:36-0400
"y'=-\dfrac{xy}{x^2+y^2}"



Substitute "y=ux"

"y'=u+xu'"


"u+xu'=-\dfrac{x^2u}{x^2+x^2u^2}"

"xu'=-\dfrac{2u+u^3}{1+u^2}"

"\dfrac{1+u^2}{2u+u^3}du=-\dfrac{dx}{x}"

Integrate


"\int \dfrac{1+u^2}{2u+u^3}du=-\int \dfrac{dx}{x}"


"\dfrac{1+u^2}{2u+u^3}=\dfrac{A}{u}+\dfrac{Bu+C}{2+u^2}"


"=\dfrac{2A+Au^2+Bu^2+Cu}{u(2+u^2)}"

"u^2: A+B=1"

"u^1:C=0"

"u^0:2A=1"


"A=\dfrac{1}{2}, B=-\dfrac{1}{2}"


"\int \dfrac{1+u^2}{2u+u^3}du=\dfrac{1}{2}\int \dfrac{du}{u}-\dfrac{1}{2}\int\dfrac{u}{2+u^2}du"

"=\dfrac{1}{2}\ln(|u|)-\dfrac{1}{4}\ln(2+u^2)+C_1"

"\dfrac{1}{2}\ln(|u|)-\dfrac{1}{4}\ln(2+u^2)=-\ln|x|+\ln C"

"\dfrac{x\sqrt{u}}{\sqrt[4]{2+u^2}}=C"


"\dfrac{x\sqrt{y}}{\sqrt[4]{2x^2+y^2}}=C"